Benoit is opening a new pizza shop and trying to estimate how much pepperoni he'll need to order. A recent randomized survey of 80 families in his area found that 29% of them ordered pepperoni on their pizza. Develop a 95% confidence statement.


The margin of error for the 95% confidence interval is 

Answer for part 1

%

Benoit can be 95% confident that between 

Answer for part 2 and coordinate 1

% and 

Answer for part 2 and coordinate 2

% of families in his area order pepperoni on their pizzas. 


1
Expert's answer
2021-04-20T02:25:34-0400

n=80, p=0.29,q=0.71


95% confidence statement is="p\pm Z_{0.025}\sqrt{\dfrac{pq}{n}}=0.29\pm 1.96\sqrt{\dfrac{0.29\times 0.71}{80}}"


"=0.29\pm 0.0994=(0.19056,0.3894)"



Margin of error "=z\times \sqrt{\dfrac{pq}{n}}=1.96\times \sqrt{\dfrac{0.29\times 0.71}{80}}=0.0994"